OpenStudy (anonymous):

A car is traveling east at 100 mph and a truck is traveling north at 80 mph. when the car is 3 miles east of an intersection and the truck is 4 miles north of the same intersection, how fast is the distance between them changing?

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

this is one of those related rates probs

OpenStudy (danjs):



OpenStudy (danjs):

It is a right triangle, and both sides are growing at different rates

OpenStudy (danjs):

write down everything that is given first... let the car distance be x, and the truck be y.. 

OpenStudy (danjs):

y = 4 mi x = 3 mi dy/dt = 80 mph dy /dt = 100 mph FIND --- dh / dt, when x and y are those values

OpenStudy (danjs):

the relationship for the three sides is the pythagorean theorem x^2 + y^2 = h^2

OpenStudy (danjs):

take the derivative of that w.r.t time (d/dt)

OpenStudy (danjs):

have to use the chain rule, implicit differentiaon 2*x*(dx/dt) + 2*y*(dy/dt) = 2*z*(dz/dt)

OpenStudy (danjs):

you can calculate z from the pythagorean theorem , for when x=3 and y=4, tha tmoment in time

OpenStudy (danjs):

so the hypotenuse is 5 put all the things in, the only thing unknown is dz/dt, how fast the hypotenuse is changing at that instance in time

OpenStudy (danjs):

2*x*(dx/dt) + 2*y*(dy/dt) = 2*z*(dz/dt) 2**3*100 + 2*4*80 = 2*5*[dz/dt]

OpenStudy (danjs):

Overall- you have a right triangle with 2 of the sides growing longer at different rates they want you to figure the info for the third side at a certain instance in time

OpenStudy (danjs):

take a derivative w.r.t time t that relates all the rates for each side, the only thing unknown is rate dz/dt

OpenStudy (danjs):

Givens - right-triangle (north and east) y = 4 mi x = 3 mi dy/dt = 80 mph dx /dt = 100 mph FIND --- dh / dt

OpenStudy (danjs):

yeah the chain rule thing, implicit differentiation is like this here x,y,and z are all functions of time t, distance traveled = rate*time

OpenStudy (danjs):

if you want to take a derivative of x w.r.t. time t, \[\large \frac{ d }{ dt }* x\] apply the chain rule \[\frac{ d }{ dt }[x] = \frac{ d }{ dx }*\frac{ dx }{ dt }[x]\]

OpenStudy (danjs):

i have to run for a bit, i will do the next one in a few

OpenStudy (anonymous):

okay thanks.

OpenStudy (danjs):

to start if you want maximum area of a cylinder , when it has a certain surface area write the formulas for the area and surface area you want to maximize the area, have to change it to one variable by solving the surface area for radius or height, and subbing that into the area A(h) or A(r) take A '(h) or A ' (r) the max is at a critical point when A '= 0

OpenStudy (danjs):

remember doing maximum and minimum values of functions, the graph changes increasing/decreasing there... first derivative is zero The area function will have max when derivative of area function is zero

OpenStudy (danjs):

solve S.A for h sub h = into the area you have A(r) area as a function of radius , take derivative and find when 0